# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning

### v1

@@ V-F2 Experiment Settings
TABLE XIV: Object Tracking: Comp. on AIR-MOT
@@
TABLE XV: Object Tracking: Comp. on AIR-HSAO
@@
Figure 14: Qualitative results for DroneVehicle-IR when using different methods
Figure 15: Qualitative results for SARData when using different methods
V-F3 Comparison With the SOTA
An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance. On unimodal datasets such as VisDrone-DET (as shown in Table VIII) and UAVDT (as shown in Table IX), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods. The experimental results further demonstrate that RingMo-Aerial, when compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Table X), and PP-PicoDet-L follows, while YOLOX-L and YOLOX-Tiny show the worst results. Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
For the multimodal dataset, DroneVehicle (as shown in Table XI), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods. Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial. This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications. Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Table XII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as YOLOX-S, Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance. Particularly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and an mAP50 of 84.4%, highlighting its superiority in SAR modality. These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery. Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments. These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields. For optical images, as shown in Figure 13, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets. In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups. Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms. RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images. As shown in Figures 14 and 15, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence. The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images. RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for aerial remote sensing.
V-G Object Tracking
V-G1 Datasets and Evaluation Metrics
For Object Tracking, following aerial dataset is employed.
  • VisDrone-MOT. The VisDrone-MOT[169] dataset consists of 96 sequences, including 5 class such as pedestrian, car, van, bus and truck. The VisDrone-MOT train and val set are used for training and test set is used for test.
  • AIR-HSAO. The AIR-HSAO dataset[168] is designed for tracking high-speed aerial objects, derived from the JiLin-1 satellite. It contains 199 videos, though fewer instances are present due to the rapid movement of these objects. The dataset spans 20 diverse regions, including urban areas, harbors, and airports, offering complex backgrounds that challenge object detection and tracking.
V-G2 Experiment Settings
In the proposed experiment, object tracking is predicated on the detection model of RingMo-Aerial on VisDrone-DET. Specifically, the detection model trained by RingMo-Aerial will undergo an additional 5 epochs of training on the VisDrone-MOT dataset, during which the backbone network parameters will be frozen to ensure consistency. The training will be conducted with a batch size of 8, distributed as 2 per GPU across 4 NVIDIA A40 GPUs. Throughout the training phase, original images from the VisDrone-MOT dataset will be processed, accompanied by multi-scale augmentation to enhance robustness. Upon completion of training on the VisDrone-MOT dataset, the detection model will be subjected to inference on the VisDrone-MOT test set using the MOT tools provided by mmdetection [148]. Subsequently, the model’s multi-object tracking performance will be quantified and analyzed.
V-G3 Comparison With the SOTA
@@
The comparison results on the VisDrone-MOT dataset are shown in Table XIII. RingMo-Aerial can achieve the best tracking results far exceeding other methods. In the AIR-MOT dataset (as shown in TableXIV), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack. Meanwhile, in the AIR-HSAO dataset (as shown in TableXV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods. Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
Figure 16: Visualization Results on AIR-MOT: Four video sequences are selected, where each target in the sequences is indicated with different colors. The category and target ID are displayed above the bounding boxes.
@@
Futhermore, the visualization results of the AIR-MOT sequences (as shown in Figure16) clearly demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios. The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions. This visualization provides strong qualitative evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in aerial remote sensing videos.
V-H ARS-Adapter
@@
TABLE XVI: Comp. between fine-tuning methods
@@
ARS-Adapter 0.41% 74.3 64.4 31.0 51.4
@@
As previously discussed, this paper proposes the ARS-Adapter, a parameter-efficient fine-tuning method tailored for ARS vision. This section focuses on comparing the performance of the ARS-Adapter with other fine-tuning methods in ARS vision applications. The comparisons are conducted using the UAVID and VisDrone-DET datasets, employing UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively. The evaluation aims to highlight the efficacy and advantages of the ARS-Adapter in enhancing ARS vision tasks compared to conventional fine-tuning techniques.
[…]
V-I Ablation Experiments
Figure 17: The ablation experiment on UAVID dataset and UperNet is used as the segmentation head. RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
TABLE XVII: Ablation Experiments
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo 71.6 61.9 28.4 45.2
RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
TABLE XVIII: Ablation Experiments with the proposed ARS-Adapter
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
The ablation experiments in this study are conducted utilizing the UperNet segmentation model on the UAVID dataset and the Cascade RCNN detection model on the Visdrone-DET dataset. The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL. This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
The comparison is shown in Figure 17 and Table XVII. The experimental findings reveal a progressive enhancement in convergence speed and performance across the three foundation models. Notably, the RingMo-Aerial model trained with both MIM and CL exhibits the swiftest convergence and attains the optimal performance. This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pretraining strategy, affirming their mutually reinforcing nature and consistent effectiveness.
In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter. Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL. Comparative results on the UAVID and VisDrone-DET datasets are presented in Table XVIII. The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
VI Conclusion
@@
The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks. Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
[…]
Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks. This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research. Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment. Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.

### v3

@@ 5.2.6 Object Tracking
TABLE XIII: Object Tracking: Comp. on AIR-MOT
@@
TABLE XIV: Object Tracking: Comp. on AIR-HSAO
@@
The comparison results on the VisDrone-MOT dataset are shown in Tab.XII. RingMo-Aerial can achieve the best tracking results, far exceeding other methods. In the AIR-MOT dataset (as shown in Tab.XIII), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack. Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XIV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods. Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
@@
Furthermore, the visualization results of the AIR-MOT sequences demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios. The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions. This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.
5.3 Ablation Experiments
5.3.1 Ablation Experiments on pre-training
Figure 13: The ablation experiment on the UAVID dataset and UperNet is used as the segmentation head. RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
TABLE XV: Ablation Experiments
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo 71.6 61.9 28.4 45.2
RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
TABLE XVI: Ablation Experiments with the proposed ARS-Adapter
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET. The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL. This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
The comparison is shown in Fig.13 and Tab.XV. The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models. Notably, the RingMo-Aerial model trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance. This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter. Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL. Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.XVI. The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
@@ 5.3.1 Ablation Experiments on pre-training
5.3.2 Performance Comparison Between ARS-Adapter and Other Fine-Tuning Methods
TABLE XVII: Comp. between fine-tuning methods
@@ 5.3.2 Performance Comparison Between ARS-Adapter and Other Fine-Tuning Methods
ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
    Although ARS-Adapter is less accurate than global fine-tuning, it significantly outperforms other fine-tuning methods when the proportion of fine-tuning parameters is almost constant.
In the context of ARS, traditional model fine-tuning presents specific challenges. These include the propensity for overfitting due to the extensive pre-training on diverse tasks, which can lead to a model that is overly tailored to the nuances of the training data, thereby compromising its applicability to new, unseen data. Overfitting is particularly problematic in ARS, where models are not only expected to generalize across various tasks but also to adapt to the unique perspectives and resolutions inherent in aerial imagery. This contrasts with ApRS, which typically operates within a more constrained perspective and resolution, reducing the risk of overfitting to specific views. The ARS-Adapter, therefore, addresses these challenges by introducing a parameter-efficient fine-tuning approach that minimizes overfitting risks while maintaining high performance, as evidenced by the experimental results.
@@
This section compares ARS-Adapter against other fine-tuning methods, using UAVID and VisDrone-DET datasets with UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively. The aim is to assess the effectiveness of ARS-Adapter compared to conventional fine-tuning techniques.
The experimental results are presented in Tab.XVII. For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer). Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment. ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency. Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
These results indicate that ARS-Adapter maintains high performance while significantly reducing the number of parameters requiring adjustment, demonstrating its efficiency in fine-tuning.
6 Conclusion
@@
The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks. Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein. Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks. This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks. In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research. Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment. Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.
Appendix A Semantic Segmentation
A.1 Datasets and Evaluation Metrics
  • UAVID. The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid, and 15 test), and each sequence contains ten pictures. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
  • UDD6. The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images. All of the images are 4K resolution. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. The training set is used for training, and the validation set is used for testing.
In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc), and average accuracy (aAcc).
A.2 Experiment Settings
The experiment of the semantic segmentation is based on segmentation [143], UperNet [144], and Mask2Former [145] are employed as the segmentation framework. The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
Figure 14: Qualitative comparison of semantic segmentation results on the UAVID dataset using different methods. The figure illustrates differences in capturing details such as vegetation, moving vehicles, and buildings, with highlighted regions emphasizing variations in detail and performance.
Appendix B Object Detection
B.1 Datasets and Evaluation Metrics
  • UAVDT. The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with an average resolution of 1080 × 540 pixels. The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
  • AerialSARData. To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside. This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection. The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels. The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
B.2 Experiment Settings
The experiment of the object detection is based on mmdetection [150], Cascade RCNN [67] and DDQ DETR [78] are employed as the detection framework. The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used). For DDQ DETR, 36 epochs were spanned with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the learning rate is decreased by 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
Figure 15: RingMo-Aerial combined with Cascade RCNN achieves superior detection accuracy compared to the original RingMo model, particularly in complex environments and small object scenarios. It demonstrates reduced false positives and improved bounding box alignment with Ground Truth, highlighting its effectiveness for aerial imagery detection.
Figure 16: Qualitative comparison of object detection results on infrared images from the DroneVehicle-IR dataset using different methods. The figure highlights differences in the detection quality of vehicles, road elements, and pedestrian areas, with emphasized regions showcasing variations in accuracy and detail among the compared methods.
Figure 17: Qualitative comparison of object detection results on the AerialSARData dataset using different methods. The figure highlights differences in detection performance for SAR images, particularly focusing on identifying objects of interest with highlighted regions showing variations in accuracy and detection quality among the methods compared.
Appendix C Change Detection
C.1 Datasets and Evaluation Metrics
  • LEVIR. The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset. It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5-meter resolution. Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels. Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases.
  • CDD. The CDD dataset is a comprehensive collection of bi-temporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels. These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches. This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs. To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
C.2 Experiment Settings
Our model adopts BIT with Swin Transformer as the backbone. The experiments use overall accuracy (OA) and F1 and IoU scores about change detection as the primary evaluation metrics. The model is trained with an AdamW optimizer with a weight decay of 0.01 and beta values equal to (0.9, 0.999). The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
Figure 18: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset. That shows the results of aerial images, where RingMo-Aerial outperforms ResNet18 and RingMo, with fewer false positives and false negatives. RingMo-Aerial closely matches the Ground Truth, demonstrating the best accuracy in change Detection. The black region represents true negative, the white region represents true positive, the green region represents false negative, the red region represents false positive.
Appendix D Scene Classification
D.1 Datasets and Evaluation Metrics
For scene classification, the AID [152] and RESISC-45 [153] datasets are employed.
  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 scenes.
In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively. In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively. The experiments use overall accuracy (OA) as a performance metric, the most commonly used performance metric in scene classification.
D.2 Experiment Settings
The experiment of the scene classification is based on mmpretrain [154], Cascade RCNN [67] is employed as the classification framework. The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs). The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
Appendix E 3D-Reconstruction
E.1 Datasets and Evaluation Metrics
For 3D-Reconstruction, the and WHU [118] datasets are employed.
Figure 19: Qualitative results on the target tracking dataset AIR-MOT using RingMo-Aerial. Four video sequences are selected, where each target is indicated with different colors. The category and target ID are displayed above the bounding boxes, demonstrating RingMo-Aerial’s ability to effectively distinguish and track individual targets.
E.2 Experiment Settings
Our work is coded based on the PyTorch framework. The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1. During the training process, our method uses the Adam optimizer[156] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001. In addition, our method is consistent with the existing multi-stage networks[119]. First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively. Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively. During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
Appendix F Object Tracking
F.1 Datasets and Evaluation Metrics
For Object Tracking, the following aerial dataset is employed.
  • VisDrone-MOT. The VisDrone-MOT[157] dataset consists of 96 sequences, including five classes such as pedestrian, car, van, bus, and truck. The VisDrone-MOT train and val set are used for training, and the test set is used for the test.
  • AIR-HSAO. The AIR-HSAO dataset[139] is designed for tracking high-speed aerial objects derived from the JiLin-1 satellite. It contains 199 videos, though fewer instances are present due to the rapid movement of these objects. The dataset spans 20 diverse regions, including urban areas, harbors, and airports, offering complex backgrounds that challenge object detection and tracking.
F.2 Experiment Settings
In the proposed experiment, object tracking is predicated on the detection model of RingMo-Aerial on VisDrone-DET. Specifically, the detection model trained by RingMo-Aerial will undergo an additional five epochs of training on the VisDrone-MOT dataset, during which the backbone network parameters will be frozen to ensure consistency. The training will be conducted with a batch size of 8, distributed as 2 per GPU across 4 NVIDIA A40 GPUs. Throughout the training phase, original images from the VisDrone-MOT dataset will be processed, and multi-scale augmentation will be accompanied to enhance robustness. Upon completion of training on the VisDrone-MOT dataset, the detection model will be subjected to inference on the VisDrone-MOT test set using the MOT tools provided by mmdetection [150]. Subsequently, the model’s multi-object tracking performance will be quantified and analyzed.
Acknowledgments
This work was supported by the National Nature Science Foundation of China under Grant 62331027 and 62301538.